OpenStudy (anonymous):

2x + y = 3 4x = 6 – 2y use addition and subtraction

OpenStudy (anonymous):

Substitute the value of y from the first equation into the second equation.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

i got x=3/4

jimthompson5910 (jim_thompson5910):

how did you get that?

OpenStudy (anonymous):

2x+y=3 4x=6-2y y=3-2x 4x=6+3-2(3-2x) 4x=6+3-6+4 4x=9-6+4 4x=3+4 x=7/4

jimthompson5910 (jim_thompson5910):

You should have this 2x + y = 3 y = 3 - 2x ------------------ 4x=6-2y 4x=6-2(3-2x) 4x=6-6+4x 4x - 4x = 6-6 0x = 0 0 = 0 So there are infinitely many solutions

OpenStudy (anonymous):

so x and y both have infinite solutions

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

ty!!!!!!!!!!!

OpenStudy (anonymous):

the answer is y=-2x+3